OpenStudy (anonymous):

A graduated cylinder with a radius of 3 cm contains water to a height of 9 cm. Yann adds 15 steel balls, each with a diameter of 2.54 cm. What is the total volume of the contents in the cylinder after the balls are added? Use 3.14 to approximate pi and round your answer to one decimal place.

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Find the volume of the cylinder, then the volume of the balls.

OpenStudy (igreen):

\(V_{cylinder} = \pi r^2 h\) \(V_{cylinder} = \pi (3)^2 (9)\)

OpenStudy (anonymous):

im still a little confused

OpenStudy (anonymous):

im gonna need more info

OpenStudy (anonymous):

\[3*3*9=81\]